Fundamentals of Financial Management (14th Edition) Chapter 12, Problem 7 Solution Not Available. Please calculate with steps. Thanks!

Scenario Analysis Huang Industries is considering a proposed project whose estimated NPV is $12 million. This estimate assumes that economic conditions will be "average". However, the CFO realizes that conditions could be better or worse, so she performed a scenario analysis and obtained these results:

Economic Scenario:               Probability of Outcome:                 NPV:

Recession                                          0.05                                ($70 million)

Below average                                   0.20                                (25 millions)

Average                                             0.50                                 12 million

Above average                                  0.20                                  20 million

Boom                                                 0.05                                  30 million

Calculate the project's expected NPV, standard deviation, and coefficient of variation.
Answer #1

Expected NPV= 0.05*(-70)+0.20*(-25)+0.50*12+0.20*20+0.05*30=$3 mn

Standard deviation= 4.73

D F G Economic Scenario Probability Probability* NPV NPV 0.05 -3.5 Recession -70 Below Average 0.2 -25 -5 Average Above Avera

C D F G H Economic Scenario Probability Probability* NPV NPV Recession 0.05 -70 -3.5 Below Average 0.2 25 -5 Average Above Av

Coefficient of variation= standard deviation/mean=4.73/3=1.578